OpenStudy (anonymous):

the force attraction(F) between two opposite electrical charges varies inversely as the square of the distance (d) between them. If F=18 when d=10 find F when D=15?

OpenStudy (anonymous):

"the force attraction(F) between two opposite electrical charges varies inversely as the square of the distance (d) between them"\[ F(d) = \frac{c}{d^2} \]

OpenStudy (anonymous):

wio, do u have any hint besides that?

OpenStudy (anonymous):

" If F=18 when d=10" What do you think this would be written as?

OpenStudy (anonymous):

18(10)=c/d^2?

OpenStudy (anonymous):

Nope. It's \[ F(10)=18 \]

OpenStudy (anonymous):

You're misunderstand function notation.

OpenStudy (anonymous):

i cannot understand it sir? :(

OpenStudy (anonymous):

Well, for example... When you have \(f(x)\), and I say \(f=6\)... you can't do this: \[ f(x) \to 6(x) \]That's not how it works.

OpenStudy (anonymous):

i know it sir :) but i cant just answer this... i dont know the answer :(

OpenStudy (anonymous):

They told you that \[ F(d) = \frac{c}{d^2} \\ F(10)= 18 \]And they want \(F(15)\). Does that make it clear?

OpenStudy (anonymous):

if u need the answer just then its F2 =F1 *d1*d1 /(d2*d2)=18 *10*10/(15*15) = 8